two angles are complementary one contains 30 degrees more than the other . find both angles
Answer by rapaljer(4551)

Let x= first angle
x+30 = second angle

x+ x+ 30= 180
2x= 150
x= 75 degrees
x+30 = 105 degrees

Check 75+105= 180

R^2